# Version dated: November 6, 2025

### v1

@@ 3 Methods II: Inference
We adopt HMC (duane_hybrid_1987), a gradient-based sampling method that is particularly well-suited for traversing high-dimensional and strongly correlated posterior distributions. By introducing auxiliary momentum variables and simulating Hamiltonian dynamics, HMC is able to propose distant, informed updates that preserve detailed balance while reducing random walk behavior (neal_mcmc_2011). This improves exploration of the parameter space, accelerates mixing, and lowers autocorrelation compared to standard MCMC methods.
@@
Each HMC update requires repeated computation of the likelihood gradient. In partially observed CTMCs, the likelihood depends on matrix exponentials of the infinitesimal rates. Differentiating these matrix exponentials is computationally demanding, typically requiring at least ${\cal O}(K^{3})$ operations per evaluation. Since the number of distinct rates scales quadratically with the number of states, the total gradient costs at least ${\cal O}(K^{5})$ per iteration. For large state spaces, this makes a naíve implementation of HMC infeasible.
@@
li_neural_2019 demonstrate that replacing the exact gradient of the log-posterior with an approximation thereof preserves the essential properties of HMC—such as reversibility and volume preservation—and that the application of a Metropolis–Hastings correction step ensures the exact posterior remains the stationary distribution of the Markov chain. Building on this insight, we adopt the scalable approximation to the matrix exponential derivative proposed by magee_random-effects_2024, reducing the computational complexity of the log posterior gradient evaluation by three orders of magnitude, making it quadratic in the number of states $K$ with respect to all rate parameters simultaneously. In the next section, we shall provide closed-form expressions for the likelihood gradient, present the aforementioned approximation, and highlight its crucial role.
@@ 3.1 Likelihood’s gradient computations
The difficulty of computing the likelihood gradient with respect to the transformed (un-normalized) rates $\theta_{ij}=g^{-1}(q_{ij})$ varies depending on the sampling scenario. For fully observed CTMCs, the gradient is straightforward from Equation 2, so it is omitted here. In contrast, for partially observed CTMCs, $p\left(\mathbf{y}\right)$ involves multiple matrix exponentials that must be differentiated. Section 3.1.1 reviews a benchmark method for computing these exact gradients and shows that its cost quickly becomes prohibitive as the state space grows. Section 3.1.2 then demonstrates how adopting an approximation of those gradients is crucial in reducing that cost and in making inference feasible, while Section 3.1.3 compares our approach with exact gradient computations based on automatic differentiation. We also apply the approximation directly to sequentially observed CTMCs with known sampling times, where the finite-time transition probability matrices do not depend on the normalizing constant. Conversely, for the other partially observed processes, Section 3.1.4 provides an explicit derivation of the gradient of $p\left(\mathbf{y}\right)$ with respect to the transformed un-normalized rates accounting for the presence of the normalization constant.
@@ 3.1.1 Gradient of the matrix exponential: Costly exact method
najfeld_derivatives_1995 review several strategies for computing the gradient of a matrix exponential. Among these, one of the most popular and straightforward approaches relies on an augmentation method. Specifically, to compute the gradient of $e^{t\mathbf{\Lambda}}$ with respect to $\lambda_{ij}$, we construct the block matrix
@@
where $\mathbf{E}_{ij}$ denotes a matrix of zeros except for a single one in position $ij$. Exponentiating $\mathbf{B}$ and extracting its upper-right block yields the Fréchet derivative of $e^{t\mathbf{\Lambda}}$ in the direction of $\mathbf{E}_{ij}$—that is, the gradient with respect to $\lambda_{ij}$. The computational cost of this exponentiation scales cubically with twice the dimension of $\mathbf{\Lambda}$. Alternative analytical methods exhibit comparable asymptotic complexity, differing primarily by constant factors. For instance, kalbfleisch_analysis_1985 derive a closed-form expression for the gradient when $\mathbf{\Lambda}$ is diagonalizable, which remains cubic in the number of states. However, in our setting, diagonalizability constitutes an overly restrictive assumption. To conclude, since gradients must be evaluated for every non-diagonal element of $\mathbf{\Lambda}$, the overall computational complexity of any such method is ${\cal O}(K^{5})$, rendering these approaches computationally prohibitive even for moderate $K$.
@@ 3.1.2 Gradient of the matrix exponential: Scalable approximation
An alternative way to express the exact gradient of a matrix exponential is through the following series (najfeld_derivatives_1995):
@@
where $\mathbf{E}_{ij}$ is a matrix with zero at all entries but at $ij$ where it is equal to one. Moreover, $\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{k}\right\}$ is defined recursively as $\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{0}\right\}=\mathbf{E}_{ij}$ and $\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{k}\right\}=\Big[\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{k-1}\right\},\mathbf{\Lambda}\Big]$, where $\cdot$$\cdot$] is the matrix commutator. The operational implementation of Equation 17 requires truncating the infinite series to a finite number of terms. Fortunately, didier_surprising_2024 demonstrate that retaining only the first term of the series already provides a remarkably accurate first-order approximation:
@@
When observations are sequential, the likelihood can be written as a product of the finite-time transition probabilities that the CTMC starting at a state $y_{u-1}$ ends up on state $y_{u}$ after time $t_{u}$, that is, $P_{y_{u-1},y_{u}}(t_{u})$ for each observation $u$. We can then rewrite each of those factors as an inner product as follows:
$P_{y_{u-1},y_{u}}(t_{u})=\mathbf{u}_{u-1}^{\prime}{\mathbf{P}}\hskip-1.99997pt\left(t_{u}\right)\mathbf{u}_{u}=\mathbf{u}_{u-1}^{\prime}e^{t_{u}\mathbf{\Lambda}}\mathbf{u}_{u},$ (19)
where $\mathbf{u}_{u}$ is a $K$-dimensional vector with all zero entries but for the one corresponding to the value of $y_{u}$. The gradient of $P_{y_{u-1},y_{u}}(t_{u})$ with respect to one rate using the aforementioned approximation is then equal to:
$\frac{\partial P_{y_{u-1},y_{u}}(t_{u})}{\partial\lambda_{ij}}=t_{u}\mathbf{u}_{u-1}^{\prime}e^{t_{u}\mathbf{\Lambda}}\mathbf{E}_{ij}\mathbf{u}_{u}.$ (20)
@@
3.1.3 A Comparison with automatic differentiation
Automatic differentiation (AD) methods have become a cornerstone of modern scientific computing, particularly in fields like machine learning and optimization. By providing efficient and exact ways to compute derivatives, they have enabled the development of complex models that were previously intractable (margossian_review_2019). In general, AD algorithms are divided into two main modes: forward mode and reverse mode. Forward-mode AD is most efficient when the function of interest depends on a small number of inputs but produces many outputs, whereas reverse-mode AD is preferable when there are many input variables and few output quantities (baydin_automatic_2018). Given our framework’s structure—with many input variables (the rates) and a single scalar output (the joint distribution)—reverse-mode AD constitutes the natural choice for gradient evaluation. This method requires a forward pass to compute the function value, which scales as ${\cal O}(K^{3})$, followed by a single backward pass to obtain all parameter gradients simultaneously. The backward pass requires computing the adjoint of the derivative of the matrix exponential, an operation of comparable complexity, also ${\cal O}(K^{3})$.
Following the same reasoning as in the previous section, within the context of CTMCs, evaluating the likelihood gradient requires at least N1N-1 adjoint computations, resulting in an overall computational complexity of ${\cal O}(NK^{3})$. Although this represents a substantial improvement over performing a full sequence of exact gradient evaluations for each individual rate parameter, it remains an order of magnitude more expensive than directly applying the proposed approximate gradient. The gap in efficiency becomes even more pronounced when considering the large $K$-invariant constant factor embedded in the ${\cal O}(NK^{3})$ cost. Specifically, computing the adjoint using the exact gradient from Equation 16 requires the evaluation of a matrix exponential of dimension $2K\times 2K$, whose cubic scaling implies a computational burden proportional to $(2K)^{3}=8K^{3}$. Moreover, even the most efficient algorithms for computing a matrix exponential require work equivalent to about $29$ matrix-matrix multiplications (najfeld_derivatives_1995).
3.1.4 Gradient with respect to the auxiliary parameters $\theta_{ij}$
In the context of partially observed CTMCs where a normalization constant is necessary for identifiability reasons, the likelihood $p\left(\mathbf{y}\right)$ depends directly only on the normalized rates $\lambda_{ij}$. Therefore, differentiating $p\left(\mathbf{y}\right)$ with respect to a transformed un-normalized rate $\theta_{ij}=g^{-1}(q_{ij})$ needs to account for its effects on each of the normalized rates (and so on the normalization constant). We now derive this gradient and evaluate it when the link function $g$ is exponential.
Since $p\left(\mathbf{y}\right)$ depends on $\theta_{ij}$ (and so on $q_{ij}$) through the normalized rates $\lambda_{ij}$, we need to apply the chain rule on $p\left(\mathbf{y}\right)$, separating the effect of $\theta_{ij}$ on each $\lambda_{ij}$, that is:
$=\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}\frac{1}{\psi}\frac{\partial q_{uv}}{\partial\theta_{ij}}-\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}q_{uv}\frac{1}{\psi^{2}}\frac{\partial\psi}{\partial\theta_{ij}}.$ (21)
Noticing that $\frac{\partial q_{uv}}{\partial\theta_{ij}}=0$ for all $uv\neq ij\text{ or }ii$, we can re-write the first summation in Equation 21 as:
$\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}\frac{1}{\psi}\frac{\partial q_{uv}}{\partial\theta_{ij}}$ $=\left(\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}-\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ii}}\right)\frac{1}{\psi}\frac{\partial q_{ij}}{\partial\theta_{ij}}=\Delta\hskip-2.5pt\left(\mathbf{y}\right)_{ij}\frac{1}{\psi}\frac{\partial q_{ij}}{\partial\theta_{ij}}$ (22)
where $\Delta\hskip-2.5pt\left(\mathbf{y}\right)_{ij}:=\left(\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}-\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ii}}\right)$ is the difference between the partial derivative with respect to a non-diagonal element and its corresponding diagonal element’s. We can then plug in Equation 22 into Equation 21 and input the exponential link function $q_{ij}=e^{\theta_{ij}}$ such that:
$\frac{\partial p\left(\mathbf{y}\right)}{\partial\theta_{ij}}$ $=\left[\Delta\hskip-2.5pt\left(\mathbf{y}\right)_{ij}-\left(\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}\lambda_{uv}\right)\pi_{u}\right]\lambda_{ij}.$ (23)
Now, notice that these expressions only depend on the distribution vector $\boldsymbol{\pi}$, the normalized rates $\lambda_{ij}$, and the gradient of $p\left(\mathbf{y}\right)$ with respect to each normalized rate. We can therefore evaluate it directly plugging in the matrix exponential approximations that we derived above.
@@ 4.1 Introduction to Bayesian phylogenetics
The phylogeny
[…]
  • $\tau=(V,E)$ is an acyclic and connected graph, with $V$ being the set of nodes and $E\subseteq V\times V$ being the set of edges. We shall refer to $\tau$ as the topology of the tree and to the edges as branches. The number of edges that are connected to a specific node is named degree of the node.
  • $\mathcal{B}$ is the set of branch lengths, a measure of the distance between two nodes directly connected by one edge.
[…]
The data generating process
The data-generating process is modeled as a CTMC that originates at the root and evolves along the tree in accordance with its conditional independence structure. We denote the realized values of the CTMC at node $i$ with $y_{i}$. Such realized values are assumed to be unknown for the internal nodes, while they are observed at the tips of the tree. We refer to such observed values with the vector $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$, where $1,2,\dots,N$ index the $N$ tips of the tree.
Phylogenetics inference
[…]
@@
where ${p}\hskip-1.00006pt\left(\mathbf{y}\hskip-1.99997pt\mid\hskip-1.00006pt{\cal F},\mathbf{Q}(\boldsymbol{\theta})\right)$ is usually referred to as the tree data likelihood. Computing the tree data likelihood implies integrating out all the values of the unobserved nodes and will be the subject of the next section.
@@ 4.2 Computing the tree data likelihood
Let $\mathbf{y}=(y_{1},\dots,y_{N})$ be the data observed at the tips of an $N$-taxa tree and $v$ be an internal node. Denote by $y_{\lfloor v\rfloor}$ the subset of $\mathbf{y}$ collecting all the descendants of $v$, while $y_{\lceil v\rceil}$ collects the non-descendants in $\mathbf{y}$. The joint distribution of the observed tip values, known as the tree data likelihood, is given by ${p}\hskip-1.00006pt\left(\mathbf{y}\hskip-1.99997pt\mid\hskip-1.00006pt{\cal F},\mathbf{Q}\right)$. A convenient method to compute this likelihood is shown in the following expression, where we omit the explicit dependence on the parameters for simplicity:
@@
where $\mathbf{p}_{v}$ is the vector of joint distributions of the descendants taxa of node $v$ conditioned on each of the values that $y_{v}$ can take, i.e., $s_{1},s_{2},\dots,s_{K}$. In contrast, $\mathbf{q}_{v}$ is the joint distribution of the non-descendants of node $v$ joint with each of the possible values that $y_{v}$ can take. In phylogenetics, $\mathbf{q}_{v}$ is named a pre-order partial likelihood vector, while $\mathbf{p}_{v}$ is referred to as a post-order partial likelihood vector.
4.2.1 Iterative algorithm for the pre- and post-order partial likelihood computations
Consider three nodes where $v$ is the parent, and $u_{1}$ and $u_{2}$ are the two children. The post-order partial likelihood vector $\mathbf{p}_{v}$ can be computed using a pruning algorithm from the tips to the root (post-order traversal) using:
$\mathbf{p}_{v}={\mathbf{P}}\hskip-1.99997pt\left(t_{u_{1}}\right)\mathbf{p}_{u_{1}}\circ{\mathbf{P}}\hskip-1.99997pt\left(t_{u_{2}}\right)\mathbf{p}_{u_{2}},$ (26)
where $\circ$ represents the Hadamard or element-wise product. In other words, we take the post-order partial likelihood vectors of the two children nodes, and we “move” them up the branch pre-multiplying them by the finite transition matrix; then, since we want both independent conditions on the children nodes to hold, we multiply them together. On the other hand, the pre-order partial likelihood vector $\mathbf{q}_{u_{1}}$ is derived in a root-to-tips tree exploration process (a pre-order-traversal) following the relation:
$\mathbf{q}_{u_{1}}={\mathbf{P}}\hskip-1.99997pt\left(u_{1}\right)^{\prime}\left(\mathbf{q}_{v}\circ{\mathbf{P}}\hskip-1.99997pt\left(u_{2}\right)\mathbf{p}_{u_{2}}\right).$ (27)
Notice that when the chosen reference node is a tip, the post-order partial likelihood is a Dirac measure. In particular, it assigns $1$ to the observed state and zero to all other possible states. Besides, the pre-order partial likelihood at the root node must be fixed a priori; we denote it by $\boldsymbol{\pi}_{\text{root}}$. Finally, notice that applying the formulas above iteratively means going through each node just once, that is the computation of the tree data likelihood is linear in the number of observed tips (felsenstein_evolutionary_1981).

### v2

@@ 3 Methods II: Inference
We adopt HMC (Duane et al., 1987), a gradient-based sampling method that is particularly well-suited for traversing high-dimensional and strongly correlated posterior distributions. By introducing auxiliary momentum variables and simulating Hamiltonian dynamics, HMC is able to propose distant, informed updates that preserve detailed balance while reducing random walk behavior (Neal, 2011). This improves exploration of the parameter space, accelerates mixing, and lowers autocorrelation compared to standard MCMC methods.
@@
Each HMC update requires the computation of the likelihood gradient. In partially observed CTMCs, the likelihood depends on matrix exponentials of the infinitesimal rates. Differentiating these matrix exponentials is computationally demanding, typically requiring at least ${\cal O}(K^{3})$ operations per evaluation. Since the number of distinct rates scales quadratically with the number of states, the total gradient costs at least ${\cal O}(K^{5})$ per iteration. For large state spaces, this makes a naíve implementation of HMC infeasible.
@@
Li et al. (2019) demonstrate that replacing the exact gradient of the log-posterior with an approximation thereof preserves the essential properties of HMC—such as reversibility and volume preservation—and that the application of a Metropolis–Hastings correction step ensures the exact posterior remains the stationary distribution of the Markov chain. Building on this insight, we adopt the scalable approximation to the matrix exponential derivative proposed by Magee et al. (2024), reducing the computational complexity of the log posterior gradient evaluation by three orders of magnitude, making it quadratic in the number of states $K$ with respect to all rate parameters simultaneously. In the next section, we shall provide closed-form expressions for the likelihood gradient, present the aforementioned approximation, and highlight its crucial role.
@@ 3.1 Likelihood’s gradient computations
The difficulty of computing the likelihood gradient with respect to the transformed (un-normalized) rates $\theta_{ij}=g^{-1}(q_{ij})$ varies depending on the sampling scenario. For fully observed CTMCs, the gradient is straightforward from Equation 2, so it is omitted here. In contrast, for partially observed CTMCs, $p\left(\mathbf{y}\right)$ involves multiple matrix exponentials that must be differentiated. Section 3.1.1 reviews a benchmark method for computing these exact gradients and shows that its cost quickly becomes prohibitive as the state space grows. Section 3.1.2 then demonstrates how adopting an approximation of those gradients is crucial in reducing that cost and in making inference feasible. We also apply the approximation directly to sequentially observed CTMCs with known sampling times, where the finite-time transition probability matrices do not depend on the normalizing constant. Conversely, for the other partially observed processes, Section S1 of the Supplement provides an explicit derivation of the gradient of $p\left(\mathbf{y}\right)$ with respect to the transformed un-normalized rates $\theta_{ij}$ accounting for the presence of the normalization constant. Finally, we refer to Section S2 of the Supplement for a comparison of our approach with computations based on automatic differentiation.
@@ 3.1.1 Gradient of the matrix exponential: Costly exact method
Najfeld and Havel (1995) review several strategies for computing the gradient of a matrix exponential. Among these, one of the most popular and straightforward approaches relies on an augmentation method. Specifically, to compute the gradient of $e^{t\mathbf{\Lambda}}$ with respect to $\lambda_{ij}$, we construct the block matrix
@@
where $\mathbf{E}_{ij}$ denotes a matrix of zeros except for a single one in position $ij$. Exponentiating $\mathbf{B}$ and extracting its upper-right block yields the Fréchet derivative of $e^{t\mathbf{\Lambda}}$ in the direction of $\mathbf{E}_{ij}$—that is, the gradient with respect to $\lambda_{ij}$. The computational cost of this exponentiation scales cubically with twice the dimension of $\mathbf{\Lambda}$. Alternative analytical methods exhibit comparable asymptotic complexity, differing primarily by constant factors. For instance, Kalbfleisch and Lawless (1985) derive a closed-form expression for the gradient when $\mathbf{\Lambda}$ is diagonalizable, which remains cubic in the number of states. However, in our setting, diagonalizability constitutes an overly restrictive assumption. To conclude, since gradients must be evaluated for every non-diagonal element of $\mathbf{\Lambda}$, the overall computational complexity of any such method is ${\cal O}(K^{5})$, rendering these approaches computationally prohibitive even for moderate $K$.
@@ 3.1.2 Gradient of the matrix exponential: Scalable approximation
An alternative way to express the exact gradient of a matrix exponential is through the following series (Najfeld and Havel, 1995):
@@
where $\mathbf{E}_{ij}$ is a matrix with zero at all entries but at $ij$ where it is equal to one. Moreover, $\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{k}\right\}$ is defined recursively as $\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{0}\right\}=\mathbf{E}_{ij}$ and $\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{k}\right\}=\Big[\left\{\mathbf{E}_{ij},\mathbf{\Lambda}^{k-1}\right\},\mathbf{\Lambda}\Big]$, where $\cdot$$\cdot$] is the matrix commutator. The operational implementation of Equation 13 requires truncating the infinite series to a finite number of terms. Fortunately, Didier et al. (2024) demonstrate that retaining only the first term of the series already provides a remarkably accurate first-order approximation:
@@
When observations are sequential, the likelihood can be written as a product of the finite-time transition probabilities that the CTMC starting at a state $y_{u-1}$ ends up on state $y_{u}$ after time $\Delta t_{u}=t_{u}-t_{u-1}$, that is, $P_{y_{u-1},y_{u}}(\Delta t_{u})$ for each observation $u$. We can then rewrite each of those factors as an inner product as follows:
$P_{y_{u-1},y_{u}}(\Delta t_{u})=\mathbf{u}_{u-1}^{\prime}{\mathbf{P}}\hskip-1.99997pt\left(\Delta t_{u}\right)\mathbf{u}_{u}=\mathbf{u}_{u-1}^{\prime}e^{\Delta t_{u}\mathbf{\Lambda}}\mathbf{u}_{u},$ (15)
where $\mathbf{u}_{u}$ is a $K$-dimensional vector with all zero entries but for the one corresponding to the value of $y_{u}$. The gradient of $P_{y_{u-1},y_{u}}(\Delta t_{u})$ with respect to one rate using the aforementioned approximation is then equal to:
$\frac{\partial P_{y_{u-1},y_{u}}(\Delta t_{u})}{\partial\lambda_{ij}}=\Delta t_{u}\mathbf{u}_{u-1}^{\prime}e^{\Delta t_{u}\mathbf{\Lambda}}\mathbf{E}_{ij}\mathbf{u}_{u}.$ (16)
@@ 4.1 Introduction to Bayesian phylogenetics
Phylogenetic inference relies on data with tree-structured dependencies. These dependencies are encoded by a graphical object called a phylogeny. A phylogeny is defined by the pair $\mathcal{F}=\langle\tau,\mathcal{B}\rangle$ consisting of an acyclic, connected topology $\tau=(V,E)$ and branch lengths $\mathcal{B}$, where $V$ is a set of nodes and $E\subseteq V\times V$ is a set of branches. The number of edges directly connected to a node is said to be the degree of the node. We denote by $N$ the number of tips of the tree, each of them having degree one. We further assume that the phylogeny is rooted and binary, that is, there is a root node with degree two and N2N-2 internal nodes with degree three. We denote the parent of node $u$ as $pa(u)$, its sibling as $sibl(u)$, and the branch length (time) from $pa(u)$ to $u$ as $b_{u}$. The data generating process consists of a CTMC that originates at the root and evolves along the tree in accordance with its induced conditional independent structure. We denote the realized value of the CTMC at node $i$ with $y_{i}$. Such realized values are assumed to be unknown for the internal nodes and the root, while they are observed at the tips of the tree. We refer to the observed values at the $N$ tips with the vector $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$.
Inference target
Given tip observations $\mathbf{\mathbf{y}}$, we seek the posterior distribution
@@
where ${p}\hskip-1.00006pt\left(\mathbf{y}\hskip-1.99997pt\mid\hskip-1.00006pt{\cal F},\mathbf{Q}(\boldsymbol{\theta})\right)$ is the tree data likelihood, while ${p}\hskip-1.00006pt\left({\cal F}\right)$ and ${p}\hskip-1.00006pt\left(\boldsymbol{\theta}\right)$ are the priors on the phylogeny and the CTMC parameters. Computing the tree data likelihood implies integrating out all the values of the unobserved nodes and will be the subject of the next section.
@@ 4.2 Computing the tree data likelihood
The joint distribution of observed tip values $\mathbf{y}=(y_{1},\ldots,y_{N})$, known as the tree data likelihood ${p}\hskip-1.00006pt\left(\mathbf{y}\hskip-1.99997pt\mid\hskip-1.00006pt{\cal F},\mathbf{Q}\right)$, can be computed efficiently by marginalizing over unobserved internal node states. For any internal node $v$, let $y_{\lfloor v\rfloor}$ denote tip values descended from $v$ and $y_{\lceil v\rceil}$ denote non-descended tips. A convenient method to compute this likelihood is shown in the following expression, where we omit the explicit dependence on the parameters for simplicity:
@@
where $\mathbf{p}_{v}$ is the post-order partial likelihood vector containing ${p}\hskip-1.00006pt\left(y_{\lfloor v\rfloor}\hskip-1.99997pt\mid\hskip-1.00006pty_{v}=s_{k}\right)$ for each state $s_{k}\in\mathbf{S}$, and $\mathbf{q}_{v}$ is the pre-order partial likelihood vector containing ${p}\hskip-1.00006pt\left(y_{\lceil v\rceil},y_{v}=s_{k}\right)$ for each state.
4.2.1 Iterative algorithm for partial likelihood computations
The vectors $\mathbf{p}_{v}$ and $\mathbf{q}_{v}$ are computed via tree traversals. Consider a parent node $v$ connected to its two children $u_{1}$ and $u_{2}$ by branches of lengths $b_{u_{1}}$ and $b_{u_{2}}$. The post-order partial likelihoods can be computed via tips-to-root traversal using the following formula:
$\mathbf{p}_{v}={\mathbf{P}}\hskip-1.99997pt\left(b_{u_{1}}\right)\mathbf{p}_{u_{1}}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{u_{2}}\right)\mathbf{p}_{u_{2}},$ (19)
where $\circ$ denotes element-wise multiplication. At the tips, $\mathbf{p}_{i}$ is a unit vector indicating the observed state. Besides, the algorithm for the pre-order partial likelihoods runs from the root to the tips applying:
$\mathbf{q}_{u_{1}}={\mathbf{P}}\hskip-1.99997pt\left(b_{u_{1}}\right)^{\prime}\left(\mathbf{q}_{v}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{u_{2}}\right)\mathbf{p}_{u_{2}}\right).$ (20)
At the root, $\mathbf{q}_{\text{root}}$ is fixed a priori to $\boldsymbol{\pi}_{\text{root}}$, which can be, e.g., the stationary distribution of the process. Both traversals visit each node once, yielding ${\cal O}(N)$ complexity (Felsenstein, 1981).